Present CRB, TAU, CWMJ, DON, FJG, RGT, RCORRADI
Action List
http://www.ing.iac.es/Intranet/engineering/opsindex.html
CWMJ received apologies for absence from KMD and was therefore chairing this meeting. CWMJ said he received a communication from KMD regarding “fault database statistics” and would be discussing this later in the meeting under AOB. The Chairperson welcomed RCORRADI to his first ops meeting prior to taking on a new role as INT/JKT telescope manager,
Reviewing the action list from the previous meeting several actions were completed or rescheduled: -
FJG requested that the AF2 software change required to position a fibre at the centre, co-ordinate (0,0) be rescheduled until 29/8/2002. An effort would be made to have the work complete prior to the next run.
CWMJ thought the INT worm and wheel problem could now be deleted but FJG said he preferred to wait until after 16/6/2002 when further tests would be made.
The meeting agreed the workaround needed to suppress the continuous error messages from AF2 system was adequate and this action should be removed from the list.
CWMJ said he had not been able to locate heat shrink or similar material to replace the damaged ‘elephant trunk’ passing through the INT polar axis, this cable reinforcing must be replaced without disconnecting the cables unless during a standown. CWMJ suggested that conventional heat shrink ‘boots’ might be used at the next aluminising standown.
FJG said the zero setting of the GHRIL nasmyth incremental encoder would become part of a more comprehensive list of TCS enhancements which would be completed on a priority bases, the meeting agreed there was no risk to INTEGRAL fibres as the rotation was ultimately restricted by limit switches.
CWMJ said the BEI encoder bulbs had finally arrived from the US and all the
WHT absolute encoders were now operational making zero setting unnecessary. CWMJ
said 20 lamps were ordered although 106 were actually delivered which is several
times the WHT’s consumption during its entire 15-year life! Because of this
proliferation of encoder bulbs CWMJ suggested the requested enhancement to
convert the WHT to LED optical encoders should be discontinued, a substantial
mechanical modifications would be needed to encoder gearboxes to accommodate
T&R encoders, T&R do not manufacture a ‘through shaft’ variant of their
encoders.
RGT circulated a communication he received from MFB regarding the maintenance of telescope optics. CRB expressed concern about the regularity and availability of mirror reflectivity measurements and RGT said he needed to discuss with MFB a policy. Some discussion followed about in-situ mirror washing and the usefulness of secondary/tertiary reflectivity measurements, some delegates saying that secondary/tertiary reflectivity measurements are made on telescopes at some observatories and would be a useful addition.
CWMJ said the Toshiba computer is the INT control room had been used to monitor the dialogue between the MCA and ICS. He went on to say that the monitoring program compresses data using an old DOS compression method and he had experienced difficulty decompressing these data for analysis.
Following discussion regarding the INT mirror support tuning parameters it was agreed that this action could be removed from the list.
FJG said RAB would implement a warning about the repeatability of the WFC integration time when a shutter speed below 2s is chosen; this action was rescheduled until 18/8/2002.
CWMJ and FJG recently used an INT dome position marked to check and adjust the dome encoder and thought the marks were adequate and this action to instate other marks was unnecessary.
CWMJ had discussed the question of WHT dome shutter leaks with DOUG. DOUG had
examined the two chambers near the top of the dome and said there was no
indication of water ingress.
WHT (CRB)
CRB reported 6 hours downtime during the 5-week period. 2 hours were lost following a NAOMI control system crash during scheduled observing time, CRB said staff at Durham University were investigating.
Two further hours were lost when stars could not be seen in the UES slit-view display. CWMJ explained that a coherent fibre bundle had been inadequately clamped and was probably disturbed when investigating a mobile (x,y direct) probe focussing problem. Subsequent to the meeting the fibre clamping problem has been addressed sky tested to ensure the guiding corrections are applied to the appropriate axis.
CRB reported an hour lost when the ‘optserver’ computer went down. CRB was concerned about the difficulty support staff have with identifying the physical whereabouts of ING computers and disk drives generally when a problem arises, DON agreed to label all fixed ING machines (Action DON).
CRB said ISIS light tightness had been greatly improved. CWMJ said he thought sponge had been placed outside the sprung collar. Subsequent to the meeting CWMJ discussed the matter with SRB who said the sponge was removed during the Worswick visit so it is assumed that the original collars are now effective.
CRB asked if the oxygen alarm had been removed from GHRIL, CWMJ said he thought TGREGORY or DOUG following a spate of false alarms removed it several months ago. CWMJ added that the Dewer is presently kept outside GHRIL on the edge of the rotating platform; CWMJ said he was concerned that the Dewer could easily become caught on the wire fence if not carefully stowed after use.
At the end of CRB’s report some members were concerned that on this occasion the report didn't include a ‘forward look’ as was normally the case. CRB outlined several instrument changes occurring in the coming weeks.
INT(TAU)
TAU reported 5:24 downtime during the 5-week period. One hour had been claimed because the clock keeping ‘mean solar time’ had been set to daylight saving time rather than UTC and an observation was missed! Some discussion followed regarding the usefulness of control room clocks. CWMJ said there were several issues; it was clear this particular confusion resulted from the clock being manually set to daylight saving time, but it was accepted that the stability of the crystal providing the time interval was not particularly good. CWMJ suggested that only the TCS time should be trusted for serious work but some members complained that the font size was too small on the display. FJG agreed to look into displaying the UTC and LST fields with larger fonts. (Action FJG).
TAU reported a small amount of time lost because the WFC filter wheel was used with an empty filter holder. CWMJ said it appeared this probably results from the wheel being out of balance causing it to run at a slightly different speed that essentially distorts the ‘time base’ over which the identification code is referenced. To solve the problem we should prepare an empty trimmed to weigh the same as a typical filter. (Action RJP, SRB)
1 hour was lost when the TCS appeared to simply crash. FJG indicated that this rare problem probably resulting from the CAMAC driver but he said that Hytec are reluctant to investigate this until we use the latest version of VMS. He went on to say various patches to VMS were available.
Much of the remaining lost time resulted from communication failures between the ICS Sparc and MCA. CWMJ said he had recently looked through fault reports going back to 14,000 and believed this type of problem was now a major component of ING downtime. RGT suggested CWMJ reported these findings in more detail at the next ops meeting (Action CWMJ).
JKT(TAU)
8:01 were lost to technical downtime during the period. One minute resulted from an apparent shutter failure during a multrun for which the meeting was unable to offer a satisfactory explanation.
Of the remainder 7 hours resulted from a filter wheel vignetting the CCD, regrettably this problem was first thought to be a CCD problem and SMT was phoned at home for advice, before the night was abandoned.
A further hour was lost to a recurring pointing problem for which nobody at the meeting could offer an explanation. Since the system worked well the following night CWMJ suggested that the TCS is restarted when this type of problem appears persistent
AOB
CWMJ said he received a communication from KMD which included statistics derived from the fault database. In this communication KMD was concerned about the increasing amount of time lost to mechanical problems. CWMJ said he had looked into this and discovered that 7 hours resulted from most of a night being lost to the vignetting filter wheel discussed at this meeting. Earlier this year considerable INT time was lost when a large volume of oil became trapped in part of the polar disk stiffening. It is understood that this produced out of balance torques dependent upon the ‘angular history’ of the HA axis and was difficult to track down.
KMD also asked what kind of information meeting delegates would like to see included in a routine analysis. The meeting felt the information presented was a good overview but a similar breakdown per telescope would be useful.
Earlier in the meeting TAU asked about TCS support following RR departure. FJG said he hoped Marion Fisher could be contracted for many of the work packages which did not need frequent access to telescopes and he would provide the on island support.
RGT suggested that this meeting should reconsider the amount of time spent discussing each telescope. CWMJ reminded the meeting that a new INT/JKT manager has recently been nominated who could presumably be expected to promote the concerns of INT/JKT users.
The meeting agreed that an effort should be made to meet monthly and CWMJ
suggested the 6/8/2002 for the next ops meeting.
Clive Jackman
Chairperson